OpenStudy (anonymous):

You can model the average national teachers’ salaries using the equation y = 9.25(1.06)n, where y is the salary in thousands of dollars and n is the number of years since 1970. What salary can a teacher expect in the year 2020?

hero (hero):

Hint: n = 2020 - 1970

OpenStudy (anonymous):

so i would plug it in like this 9.25(106)^50? @Hero

hero (hero):

Is this the original equation: \(y = 9.25(1.06)^n\)

OpenStudy (anonymous):

yes

hero (hero):

If so, then: \(y = 9.25(1.06)^{50}\) is correct

OpenStudy (anonymous):

alright thank you!

hero (hero):

Well, but something seems wrong

hero (hero):

What did you get for \(y\) after calculation?

OpenStudy (anonymous):

i got 170.386427

hero (hero):

That's what I got as well. I don't think the average teacher salary is \($170,000\) per year. Teachers don't make nearly as much.

OpenStudy (anonymous):

i dont think its an actual fact though

hero (hero):

Are you sure the equation is correct to begin with? Double check all the decimal points and the numbers

hero (hero):

It doesn't have to be an actual fact, however, it should be realistic

OpenStudy (anonymous):

one of the answer choices is 170,400

hero (hero):

I see...

hero (hero):

In that case....it seems you have found the correct choice.

OpenStudy (anonymous):

sweet! thank you for your help!

hero (hero):

I think a teacher created this question. That must be what he or she wishes their salary to be.